# A Disproof of Large Language Model Consciousness: The Necessity of Continual Learning for Consciousness

### 12 Jan 2026 v2

@@ 5 The Necessity of Continual Lenient Dependency
Proposition 5.1 (Continuity Requirement for Lenient Dependency).
@@ 5.1 Continual Learning as Continual Lenient Dependency
Lenient dependency requires that a theory’s predictions not be a priori falsified by known universal substitutions but is also not strictly dependent, and we now know from Proposition 5.1 that this must continually hold or else a theory of consciousness collapses. Here, I show that a theory involving continual learning offers such a hypothesis.
Proposition 5.3 (Invalidity of Non-learning Substitutions for Learning Systems).
Non-learning systems cannot be valid universal substitutes for systems that learn. Any purported non-learning substitute must either:
@@
    Contain internal memory: the static universal substitute may track history internally to match any given learning system’s input/output function. But by definition, any system with internal states that itself updates based on inputs, and so change its overall input/output function, is also itself a learning system. Therefore, this does not constitute a non-learning substitution.
@@
    Contain external memory: requiring input $(x,\text{history})$ rather than $x$ alone, like a large (or even infinite) static lookup table. However, this requires that the input/output is not preserved in the universal substitution, since the input is substantially different. It is not, e.g., the same input information but in a different temporal sequence; in cases of input $(x,\text{history})$ there can be orders of magnitude greater information in the substituted system’s input, thus making it a false universal substitution over the input domain.
  3. (c)
@@
    Be derivative: One might ask if a learning system could be substituted for a series of static systems instead, each of which is like a static snapshot of the system. This series might functionally mimic learning. However, testing in such a system cannot be a swap, but rather, an ongoing reconstruction. This means that any such series cannot be tested in isolation, i.e., it cannot be truly substituted in for the original, as it depends on monitoring changes in the original.
@@
Already, [59] has shown a specific case of Proposition 5.3, that plastic RNNs cannot be unfolded into an equivalent static FNN. However, this argument is not a repetition of their argument: by focusing on learning rather than on plasticity, and substitutions in general rather than just unfolded networks, it offers a broader case. E.g., in [59] many of the proofs are based on resource constraints, and do not rule out cases like an infinite static system that contain all possible histories acting as a viable substitution (which are physically impossible, but still potentially relevant to reasoning about falsification). However, Proposition 5.3 does, since it shows that input $(x,\text{history})$ violates the substitution. Note that this applies to, e.g., LLMs acting as substitutes for humans: for an LLM, the context window is fed back into the input, thus, LLMs contain external memory (Proposition 5.3(b)) and their inputs are actually quite unlike humans even if they were to act as substitutes in, e.g., a Turing Test [71] (each query actually involves the entire previous conversation fed into a static system).
Given Proposition 5.3’s truth, the failure of all static systems to offer viable universal substitutions for learning systems means that the set of universal substitutions is likely a smaller subset when it comes to theories of consciousness that involve learning. In other words, theories of consciousness that involve learning are harder to a priori falsify by finding universal substitutions. Consider how much more it requires than matching static input/output functions: A universal substitution for a learning system needs to somehow match the entire history of the learning process, while still at the same time ensuring equivalency across input/output! This is not to say that these results currently show that all universal substitutions for learning systems are impossible; the clearest option seems to be, e.g., simulations of a system on a computer, which might, depending on the theory, create pathological mismatches. Addressing this is a direction for future research. However, they do rule out the entire class of static substitutions for theories of consciousness within which predictions are grounded in or require continual learning.
@@
Finally, it is worth noting that there exists an obvious relationship regarding the overflow of predictions with regards to inferences to the classic philosophical issues of phenomenological consciousness vs. access consciousness [17]. Learning is about counterfactuals, generalization, and dispositions, which all involve information far beyond any given particular test. In other words, a theory’s predictions grounded in learning (say, focused on physical plasticity) “overflow” the access of individual experimental inferences, since the subject of those predictions (like plastic changes) implicitly have effects on a much larger domain (the entire set of all possible experiments, not just the current experiment).
@@ 6.1 A Progressive Research Program on Consciousness
The literature around formal requirements theories of consciousness, which so far has focused on falsification [31] and non-triviality [45] and also “toy” models [4], may to outsiders at first seem a small subfield of the neuroscience of consciousness focused on highly-theoretical “armchair” arguments. For this reason, some in the literature have argued that formal or theory-based arguments should take a backseat to empirical research, and we should not put strong constraints on theories of consciousness. [72]
@@
It seems likely that some will argue against these results by debating the definition of falsification, such as pointing out that while a simpler Popperian definition is used herein [62], in comparison Lakatos offers a more complex form of falsification [55]. However, even if some more complicated conception of falsification complicates the Kleiner-Hoel dilemma (we have no reason to think it would rule it out, given that such alternatives are often just “refined” versions of Popperian forms of falsification), it would still be a move of last resort: if we have clear and simple falsification criteria for a science of consciousness, and we know certain theories do pass these without cases of pathological mismatches for identical input/output systems, and others do not, we should be very careful about relaxing the standards of science until it’s shown to be impossible to make progress without doing so.
The point of this paper goes beyond its contents: the goal is to illustrate how a progressive approach to a science of consciousness might take place via painstakingly working out all the requirements theories of consciousness need to satisfy in order not to be a priori falsified and yet still remain non-trivial and scientifically informative. That this already rules out a great number of contemporary theories, gives us the structure of a strong disproof of LLM consciousness, and points us toward the capability of continual learning, is the kind of actually surprising information that indicates a progressive research program.
Indeed, this allows us to define a precise target for theoretical consciousness research: a Minimal Non-trivial and Falsifiable Theory of Consciousness (MNFTC). An MNFTC is a theory of consciousness that, via satisfying continual lenient dependency, does not succumb to any of the logical problems around falsifiability and testability that the super-majority of contemporary theories fall into. It seems that somehow equating learning with consciousness in some way may be the clearest path to this, although defining such a theory is left for future research.
6.2 Constraints on Theories Rule out the Consciousness of Many Artificial Systems (But Not All)
It would be easy to misunderstand this paper as a restatement of biological naturalism or some other theory that gives human consciousness undue priority. This is very much the opposite: here, I take seriously the reports of non-human systems as containing information about a theory of consciousness. The results are that:
  • Theories of consciousness in LLMs are highly constrained due to their proximity (via transformation or substitution) to systems for which there are no non-trivial theories of consciousness possible. Therefore, LLMs are not conscious.
  • The “lenient dependency” required for theories of consciousness to escape issues of falsification (or unfalsifiability) must be continually true and present in falsifiable and non-trivial theories.
  • True continual learning appears to escape these issues of falsification in humans. This implies that continual learning, widely known as the Holy Grail of future AI success, is intrinsically linked to consciousness, and could be implemented in non-biological systems.
One potential limitation is simply that these claims are too strong, indicating that there may be flaws in the reasoning here (if only due to my poor instantiation of it). Additionally, there is a large amount of future research to be done sketching out more formally what this progressive research program would consist of, delineating its boundaries, and also evaluating potential theories that actually fit those constraints, and connecting this to other work [50]. Therefore, it’s worth noting there may be other cases of lenient dependency beyond continual learning that are more exotic (speculative alternatives include, say, observer effects in quantum mechanics being somehow immune to substitutions [36]). However, learning has a strong claim by being already reasonable to include in inferences about consciousness: could there really be experiences that leave no trace whatsoever?
@@ 6.3 Continual Learning as Central to Consciousness
It is not a coincidence that LLMs, despite their intelligence, still feel like “ghosts” [43]. If consciousness is truly based on continual learning (or requires it for some further property), then it makes sense that the substitution of a frozen input/output echo of humans would be technically impressive but fail as our substitutions in ways that are hard to specify or quantify. Often science proceeds spurred on by metaphors from technology: LLMs being close substitutions to humans for many input/output text-based or screen-based tasks, but then also being somehow incomplete substitutions in ways difficult to quantify or specify, seems to preface the results herein about how input/output equivalence is a less valid target than first appears.
@@
Of course, it is worth noting that there are some theories of consciousness that are already focused on learning and plasticity [25, 60]. The results herein should make these prime candidates for exploration, adoption, and testing. It is a matter of future research to what degree existing theories explicitly fit the requirements of continual lenient dependency, as some may be based only on past learning, rather than current learning.
Intriguingly, the disproof herein does not rule out LLM consciousness during their training in the same way it does for the deployed “dead” artificial neural network; however, nor does it guarantee their consciousness in any way (i.e., training might not be similar to the actual kind of continual learning involved in consciousness. nor be technically continual in practice), and so this remains an area reserved for future research. When it comes to training, it should be noted that there are detectable differences between how humans vs. LLMs learn, already indicating that the learning process itself may differ significantly and indicating that LLM consciousness may not hold during training [61].
Finally, with regard to humans, a view of the biological brain as being a gestalt continuously updated by every experience is, at least, scientifically evocative, and not contradicted by the neuroscience of learning. E.g., place fields can develop in a single trial [15], and significant plasticity can occur over seconds [24]; even the Overfitted Brain Hypothesis hypothesizes that dreams may have evolved to combat daily overfitting [37]. It may indeed be that every experience must leave a trace for it to be an experience at all.
@@ 7 Acknowledgments
Thank you to the supporters of Bicameral Labs; thanks as well to Johannes Kleiner for conversations and feedback.

### 19 Jan 2026 v3

@@ 5 Continual Lenient Dependency
Proposition 5.1 (Continual Requirement for Lenient Dependency).
@@ 5.1 Continual Learning as Continual Lenient Dependency
Lenient dependency requires that a theory’s predictions not be a priori falsified by known universal substitutions but is also not strictly dependent, and we now know from Proposition 5.1 that this must continually hold or else a theory of consciousness collapses. Here, I show that a theory involving continual learning offers such a hypothesis by significantly complicating the viability of universal substitutions.
Naturally, a theory of consciousness involving learning would involve testing learning rather than merely static I/O (including evidence of learning in inferences). Specifically, we can define for some static function f, there is $\Delta f$, which here represents the change from learning.
Definition 5.3 (Static System).
For a system $s$ with an I/O function, $f$, then $\Delta f=0$ if identical-in-content inputs performed at different times cannot possibly yield different outputs, excluding randomness or internal changes not caused by inputs.
Proposition 5.4.
Static Systems (Definition 5.3), such as a lookup table $L$ implementing the function $f$, would be unable to act as a universal substitution for a learning system in a consciousness-testing regime sensitive to $\Delta f$, since for the original system $\Delta f\neq 0$ and by Definition 5.3 $L$’s $\Delta f=0$.
Corollary 5.5.
Static Systems which implement a static $f$ cannot be valid universal substitutes for systems that learn and for which $\Delta f\neq 0$. Any universal substitute must either actually itself learn as well (via things like changes to internal memory), or be a non-learning system which either:
@@
    Stores external memory. A system may input $(x,\text{history})$ rather than $x$ alone, like a static lookup table. This would violate Definition 5.3 of being a Static System, since identical inputs cannot yield differentiable outputs. Additionally, I/O will necessarily expand in scope to include another kind of data (history) rather than, e.g., sensory processing, leading to different domains over the input (and possibly output if a loop is necessary), indicating a failure to substitute.
@@
    Is experimentally derivative. A series of static substitutions might replace the trajectory of a learning system. However, testing would require ongoing reconstruction, i.e., the series cannot be truly substituted in for the original and experimented on in isolation, as the series depends on continually monitoring changes in the original, upon which the actual experiments that include $\Delta f$ must also be performed.
@@
This is supported by the literature: indeed, a specific case that a plastic recurrent neural network actually cannot be unfolded into an equivalent static FNN has been shown [75]. However, by focusing on learning rather than on plasticity, and substitutions in general rather than just unfolded networks, this offers a broader case. E.g., in [75] many of the proofs are based on resource constraints, and do not rule out cases like an infinite static system that contain all possible histories acting as a viable substitution (which are physically impossible, but still potentially relevant to reasoning about falsification). However, Corollary 5.5 indicates that even such instances are impossible. Note that this applies to, e.g., LLMs acting as substitutes for humans: for an LLM, the context window is fed back into the input, thus, via Corollary 5.5(b), baseline LLMs mimic learning via external memory. Fed the same $\text{input}(x,\text{history})$ at any time, their output probabilities would be identical. Additionally, their I/O is actually quite unlike humans even if they were to act as substitutes in, e.g., a Turing Test [89] (each query actually involves the entire previous conversation fed into a static system). This covers the case for what is often referred to as “in-context learning” for LLMs.
Given Proposition 5.4’s truth, the failure of all static systems to offer viable universal substitutions for learning systems means that the set of universal substitutions is a smaller subset when it comes to theories of consciousness that involve learning. In other words, theories of consciousness that involve learning are harder to a priori falsify by finding universal substitutions. Examples of substitutions in the literature have relied entirely on known mathematical methods like the universal approximation theorem [37, 48] to arbitrarily approximate some continuous function $f$, not some $\Delta f$. This is not to say that the results herein currently show that all universal substitutions for learning systems are impossible. However, they do rule out the class of static substitutions for theories of consciousness within which predictions are grounded in or require continual learning, and point toward testing necessarily including learning.
Learning also offers clear ways to further complexify the issue that seem to make substitutions harder to define or implement: for instance, learning involves generalization to new never-before-seen inputs, but the status of which under question for I/O substitutes—can you really substitute every possible history once learning is involved, without necessarily relying on infinities? Additionally, learning systems may not be enumerable in the same way as static ones, since enumeration of inputs and outputs leads to plastic changes over those very inputs and outputs. Indeed, focusing on learning-based theories of consciousness may connect to the hypothesis that consciousness may involve what Geoffrey Hinton dubbed “mortal computation” (for a line of argument that consciousness must be such a mortal computation, see [61])—defined as computation unable to be divorced from its substrate [45]. With regards to continual learning, there could be an analogous version of “mortal learning,” wherein changes caused by input depended on no static memory being read/written. If so, this would rule out any universal substitution that could be viewed (under some perspective) as undergoing separable read/write execution as part of the input of the system, thus changing the scope of I/O, for similar reasons as in external memory in Corollary 5.5, and thus failing to be a universal substitution for a mortal learning system.
@@
Finally, it is worth noting that there exists a relationship here to the classic philosophical issues of phenomenological consciousness vs. access consciousness [17]. Learning is about counterfactuals, generalization, and dispositions, and so a theory’s predictions grounded in learning (say, focused on physical plasticity) “overflow” the access of individual experimental inferences, since the subject of those predictions (like plastic changes) implicitly have effects on a much larger domain (the entire set of all possible experiments, not just the current experiment).
@@ 6.1 A Progressive Research Program on Consciousness
The literature around formal requirements for theories of consciousness, which so far has focused on falsification [37] and non-triviality [57] and also “toy” models [4], may to outsiders at first seem a small subfield of the neuroscience of consciousness focused on highly-theoretical “armchair” arguments. For this reason, some in the literature have argued that formal or theory-based arguments should take a backseat to empirical research, and we should not put strong constraints on theories of consciousness. [90]
@@
The point of this paper goes beyond its contents: the goal is to illustrate how a progressive approach to a science of consciousness might take place via painstakingly working out all the requirements theories of consciousness need to satisfy in order not to be a priori falsified and yet still remain non-trivial and scientifically informative. That this already rules out a great number of contemporary theories, gives us the structure of a disproof of LLM consciousness, and points us toward the capability of continual learning, is the kind of actually surprising information that indicates a progressive research program.
Indeed, this allows us to define a precise target for theoretical consciousness research: a Minimal Non-trivial and Falsifiable Theory of Consciousness (MNFTC). An MNFTC is a theory of consciousness that, via satisfying continual lenient dependency, does not succumb to any of the logical problems around falsifiability and testability that the supermajority of contemporary theories fall into. Any universal substitutions are more like simulations of the properties that a theory’s predictions are sensitive to. There is a large amount of future research to be done sketching out more formally what this progressive research program would consist of, delineating its boundaries, and also evaluating potential theories that actually fit those constraints, and connecting this to other work [63]. And while this paper focuses on learning as one option, it’s worth noting there may be other cases of lenient dependency beyond continual learning that are more exotic (speculative alternatives include, say, observer effects in quantum mechanics being somehow immune to substitutions [44]).
@@ 6.2 Continual Learning as Central to Consciousness
It is not a coincidence that LLMs, despite their intelligence, still feel like “ghosts” [53]. If consciousness is truly based on continual learning (or requires it for some further property), then it makes sense that the substitution of a frozen I/O echo of humans would be technically impressive but fail as our substitutions in ways that are hard to specify or quantify. Often science proceeds spurred on by metaphors from technology: LLMs being close substitutions to humans for many I/O text-based or screen-based tasks, but then also being somehow incomplete substitutions in ways difficult to quantify or specify, seems to preface the results herein about how I/O equivalence is a less valid target than first appears.
@@
Of course, it is worth noting that there are some theories of consciousness that are already focused on learning and plasticity [29, 76], including (depending on variant and interpretation) some popular theories based on minimizing prediction error and predictive processing that involve updating representations like world models [41, 84, 28, 80]. The results herein should make these prime candidates for exploration and testing. It is a matter of future research to what degree existing theories explicitly fit the requirements of continual lenient dependency, as some may be based only on past learning, rather than current learning.
Intriguingly, the disproof herein does not rule out LLM consciousness during their training in the same way it does for the deployed “dead” artificial neural network; however, nor does it guarantee their consciousness in any way (i.e., training might not be similar to the actual kind of continual learning involved in consciousness, nor be technically continual in practice), and so this remains an area for future research. Even if continual learning (which satisfies Proposition 5.1, the Continual Requirement for Lenient Dependency) were to be introduced in LLMs (see attempts such as, e.g., [12]), the results herein should give us pause. They indicate that non-conscious systems can produce highly elaborate claims to consciousness and this should considerably raise the bar of suspicion. Additionally, continual learning here is given as merely a necessary condition, not a sufficient one, and the results also indicate that LLM consciousness may be limited by degree of plasticity and how that plasticity is distributed (e.g., a small module that continually learns appended to an LLM would not necessarily render the LLM conscious of its conversations). At least when it comes to training, it should be noted that there are detectable differences between how humans vs. LLMs learn, already indicating that the learning process itself may differ significantly and so indicating that LLM consciousness may not hold during training [78].
Finally, with regard to humans, a view of the biological brain as a gestalt being continuously updated by literally every experience is, at least, scientifically evocative, and not contradicted by the neuroscience of learning. E.g., place fields can develop in a single trial [15], and significant plasticity can occur over seconds [26]; even the Overfitted Brain Hypothesis hypothesizes that dreams may have evolved to combat daily overfitting [46]. It may indeed be that every experience must leave a trace for it to be an experience at all.
@@ 7 Acknowledgments
Thank you to the supporters of Bicameral Labs; thanks as well to Johannes Kleiner and Abel Jansma for conversations and feedback.